OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned?

OpenStudy (anonymous):

I got 52 women for the answer.

OpenStudy (anonymous):

the options are only 44, 40, 76, and 80

OpenStudy (anonymous):

I'll try again

OpenStudy (anonymous):

okay thanks !

OpenStudy (anonymous):

The way I'm setting it up is: \[y + x = 120\] \[3y - 2x = 20\] where x = men and y = women. Multiply first equation by 2. \[2y + 2x = 240\] \[3y - 2x = 20\] Simplify. \[5y = 260\] Divide both sides by 5. \[y = 52\] And as we said earlier, y = women. \[women = 52\]